$SPAD/src/input fname1.input

Timothy Daly December 3, 2016

Abstract

Contents

2

1 License

— license —

--Copyright The Numerical Algorithms Group Limited 1991.

———-

— * —

)set break resume
)spool fname1.output
)set message test on
)set message auto off
)clear all
--S 1 of 18
fn: FileName
--R
--R                                                                   Type: Void
--E 1
--S 2 of 18
fn := "/spad/src/input/fname.input"
--R
--R
--R   (2)  "/spad/src/input/fname.input"
--R                                                               Type: FileName
--E 2
--S 3 of 18
directory fn
--R
--R
--R   (3)  "/spad/src/input"
--R                                                                 Type: String
--E 3
--S 4 of 18
name fn
--R
--R
--R   (4)  "fname"
--R                                                                 Type: String
--E 4
--S 5 of 18
extension fn
--R
--R

--R   (5)  "input"
--R                                                                 Type: String
--E 5
--S 6 of 18
fn := filename("/u/smwatt/work", "fname", "input")
--R
--R
--R   (6)  "/u/smwatt/work/fname.input"
--R                                                               Type: FileName
--E 6
--S 7 of 18
objdir := "/tmp"
--R
--R
--R   (7)  "/tmp"
--R                                                                 Type: String
--E 7
--S 8 of 18
fn := filename(objdir, "table", "spad")
--R
--R
--R   (8)  "/tmp/table.spad"
--R                                                               Type: FileName
--E 8
--S 9 of 18
fn := filename("", "letter", "")
--R
--R
--R   (9)  "letter"
--R                                                               Type: FileName
--E 9
--S 10 of 18
exists? "/etc/passwd"
--R
--R
--R   (10)  true
--R                                                                Type: Boolean
--E 10
--S 11 of 18
readable? "/etc/passwd"
--R
--R
--R   (11)  true
--R                                                                Type: Boolean

--E 11
--S 12 of 18
readable? "/etc/security/passwd"
--R
--R
--R   (12)  false
--R                                                                Type: Boolean
--E 12
--S 13 of 18
readable? "/ect/passwd"
--R
--R
--R   (13)  false
--R                                                                Type: Boolean
--E 13
--S 14 of 18
writable? "/etc/passwd"
--R
--R
--R   (14)  true
--R                                                                Type: Boolean
--E 14
--S 15 of 18
writable? "/dev/null"
--R
--R
--R   (15)  true
--R                                                                Type: Boolean
--E 15
--S 16 of 18
writable? "/etc/DoesNotExist"
--R
--R
--R   (16)  true
--R                                                                Type: Boolean
--E 16
--S 17 of 18
writable? "/tmp/DoesNotExist"
--R
--R
--R   (17)  true
--R                                                                Type: Boolean
--E 17

--S 18 of 18
fn := new(objdir, "xxx", "yy")
--R
--R
--I   (18)  "/tmp/xxx1420.yy"
--R                                                               Type: FileName
--E 18
)spool
)lisp (bye)

———-

References^{[1] nothing}

7